### v1

# ZeroPur: Succinct Training-Free Adversarial Purification

Abstract
Adversarial purification is a kind of defense technique that can defend various unseen adversarial attacks without modifying the victim classifier. Existing methods often depend on external generative models or cooperation between auxiliary functions and victim classifiers. However, retraining generative models, auxiliary functions, or victim classifiers relies on the domain of the fine-tuned dataset and is computation-consuming. In this work, we suppose that adversarial images are outliers of the natural image manifold and the purification process can be considered as returning them to this manifold. Following this assumption, we present a simple adversarial purification method without further training to purify adversarial images, called ZeroPur. ZeroPur contains two steps: given an adversarial example, Guided Shift obtains the shifted embedding of the adversarial example by the guidance of its blurred counterparts; after that, Adaptive Projection constructs a directional vector by this shifted embedding to provide momentum, projecting adversarial images onto the manifold adaptively. ZeroPur is independent of external models and requires no retraining of victim classifiers or auxiliary functions, relying solely on victim classifiers themselves to achieve purification. Extensive experiments on three datasets (CIFAR-10, CIFAR-100, and ImageNet-1K) using various classifier architectures (ResNet, WideResNet) demonstrate that our method achieves state-of-the-art robust performance. The code will be publicly available.
1 Introduction
Recent studies show that adding carefully crafted, imperceptible perturbations to natural examples can easily fool deep neural networks (DNNs) to make wrong decisions [42, 11]. This potential vulnerability underlying their remarkable performance raises a significant challenge for security-critical applications. Thus, designing efficient adversarial defense techniques is necessary for real-world applications in DNNs.
One kind of adversarial defense technique is adversarial training [28, 54, 27, 23], which involves adversarial examples in the model training, enabling the model to adapt adversarial perturbations empirically. However, these approaches typically require huge computational resources [37, 48] and suffer from performance degradation [24] in the presence of unseen attacks that are not involved in training. This limitation hinders the application of adversarial training in realistic scenarios.
Figure 1: An illustration of ZeroPur.
Different from adversarial training, adversarial purification [51, 38, 30, 29] aims to remove adversarial perturbations in adversarial examples. These methods do not require adversarial examples in the model training and effectively defend against unseen attacks, making them more applicable in real-world scenarios. However, the existing purification methods often depend on external generative models [10, 40] or cooperation between external auxiliary functions and the victim classifier [38, 29, 19]. Retraining generative models, parameterized auxiliary functions or the classifier relying on the domain of the fine-tuned dataset is computationally demanding and restricts their flexibility.
Inspired by the natural image manifold hypothesis, we suppose adversarial images are outliers of the natural image manifold and the purification process is to return them to this manifold. We present a simple adversarial purification method named ZeroPur, where "Zero" means our method does not need to train any external models, parameterized auxiliary functions, and victim classifiers. As illustrated in Fig. 1, given the adversarial image, ZeroPur comprises two stages, i.e., Guided Shift (GS) and Adaptive Projection (AP), to purify the given image. In detail, GS guides adversarial examples towards the blurred counterparts to target natural images on the manifold to obtain the shifted embeddings. This is achieved by iteratively pulling the distance between adversarial examples and a set of their blurred counterparts in the embedding space. Since the inherent limitations of the low-quality embeddings from blurred counterparts, adversarial examples may not precisely return to the target locations, but these blurred counterparts do provide a reasonable direction. After that, we introduce AP to construct the directional vectors based on these shifted embeddings which provide a reference direction of the movement of adversarial images. The adversarial images are projected adaptively onto the manifold by projection maximization in this direction. Despite its simplicity, ZeroPur consistently outperforms most state-of-the-art adversarial training and purification methods.
The main contributions of the current work are as follows:
  • We analyze the relationship between adversarial attack and adversarial purification based on the natural image manifold hypothesis, and show that a simple blurring operator can bring adversarial examples that are out of the natural image manifold closer to the manifold.
  • We present a succinct adversarial purification approach named ZeroPur including two stages: Guided Shift and Adaptive Projection, which requires no retraining of external models, parameterized functions, and the victim classifier.
  • Extensive experiments demonstrate that the proposed approach outperforms most state-of-the-art auxiliary-based adversarial purification methods and achieves competitive performance compared to other external model-based purification methods.
2 Review of Literature
Adversarial training (AT). AT [28, 54, 27, 23] improves the robustness of DNNs by integrating adversarial examples into the training data and reformulating the optimization objective. However, the computational cost of adversarial training is significantly huge due to the necessity of repeatedly performing backpropagation to craft adversarial examples. While recent works investigate reducing the time cost of adversarial training, they are still restricted by issues such as low robust performance [37, 2] and other unexpected results [2, 25] (i.e., catastrophic overfitting). Moreover, even if a model is robust against a specific set of known attacks, it is still fragile against other unseen attacks that were not involved in training [20, 24, 8].
External model-based adversarial purification (EBP). Samangouei et al. [35] propose defense-GAN, a generator that models the distribution of natural images, which enables the transformation from adversarial examples to natural images. Song et al. [39] assume that adversarial examples primarily reside in the low probability density region of the training distribution, and design PixelDefend to approximate this distribution using the PixelCNN [44]. Recently, utilizing score-based models [51] and diffusion models [45, 30] as purification models is introduced and shown to achieve significantly improved robust performance. More recently, Lin et al. [26] propose a framework called AToP that fine-tunes the purification model adversarially to integrate the benefits of both adversarial training and adversarial purification. However, these works rely on external generative models, substantial datasets (e.g., crafting adversarial examples) and computational resources.
Auxiliary-based adversarial purification (ABP). In contrast to the above two approaches, ABP tends to introduce an auxiliary function to cooperate with the classifier to purify adversarial images. Shi et al. [38] propose a lightweight purification method SOAP, which uses self-supervised loss to realize online purification. SOAP no longer depends on generative models but requires classifiers to incorporate the corresponding auxiliary loss (self-supervised loss) in the training stage. To further reduce the time cost of purification, recent works introduce parameterized auxiliary functions, allowing purification to be accomplished through the exclusive training of these functions. Such auxiliary functions are designed to be lightweight compared to classifiers so that they can significantly reduce the time cost. For example, Mao et al. [29] introduce a two-layer network that estimates the contrastive features, which can purify the adversarial images. Hwang et al. [19] propose AID-Purifier, an auxiliary discriminator based on information maximization principles that can transform adversarial images into natural images.
3 ZeroPur
3.1 Adversarial Purification in The Natural Image Manifold
Following the natural image manifold hypothesis [17], natural images are assumed to reside on a specific manifold known as the natural image manifold. Given the classification loss function $\ell$, the learning process of DNNs finding optimal parameters $\theta$ can be formulated by:
$\min_{\theta}\ell(f\circ g(\bm{x}+\bm{\delta}^{*}),y;\theta),$ (1)
[…]
Adversarial attack crafts adversarial examples by optimizing the following objective:
$\max_{\|\bm{\delta}^{*}\|\leq\epsilon}\ell(f\circ g(\bm{x}+\bm{\delta}^{*}),y),$ (2)
where $\epsilon$ is the maximal norm which defines the set of allowed perturbations for a given example $\bm{x}$. In most contexts, such as The Projected Gradient Descent (PGD) [28], $\bm{\delta}^{*}$ is approximated by the local worst-cast $\bm{\delta}$.
We are starting with formulating the forward process of $f$. Suppose $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ is twice differentiable, then the forward of $f$ can be defined as:
$f_{1...l}(\bm{x})=f_{1}\circ f_{2}\circ...\circ f_{l}(\bm{x})=f_{l}(f_{l-1}(f_{2}...f_{1}(\bm{x}))),$ (3)
where $f_{i}$ typically refers to the output of the $\textit{i}^{th}$ layer in the function $f$. Based on the Taylor series, we have the following theorem:
Theorem 1.
Suppose $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ is twice differentiable at point $\bm{x}$. Let $e_{1}=\nabla f_{1}(\bm{x})^{T}\bm{\delta}+\frac{1}{2}\bm{\delta}^{T}\nabla^{2}f_{1}(\bar{\bm{x}})\bm{\delta}$, and there exists $w_{1}\in[0,1]$ such that $\bar{\bm{x}}=w_{1}\bm{x}+(1-w_{1})(\bm{x}+\bm{\delta})$, then we have the forward of $f$:
$f_{1...l}(\bm{x}+\bm{\delta})$ $=f_{1...l}(\bm{x})+e_{l}(\bm{\delta}),$ (4)
$where\ \ \ e_{l}(\bm{\delta})=\nabla f_{l}(f_{1...l-1}(\bm{x}))^{T}$ $e_{l-1}(\bm{\delta})+\frac{1}{2}e_{l-1}(\bm{\delta})^{T}\nabla^{2}\bar{f}_{1...l-1}(\bm{x})e_{l-1}(\bm{\delta}).$ (5)
where there exists $w_{l}\in[0,1]$ such that $\bar{f}_{1...l-1}(\bm{x})=w_{l}f_{1...l-1}(\bm{x})+(1-w_{l})(f_{1...l-1}(\bm{x})+e_{l-1}(\bm{\delta}))$.
For notational simplicity, we denote $f_{1...l}(\bm{x})=f(\bm{x})$ and $e_{l}(\bm{\delta})=e(\bm{\delta})$ when $l$ is the last layer of $f$. Theorem 1 describes how the perturbation $\bm{\delta}$ is amplified layer by layer during the forward of $f$ so that the embedding of natural image $\bm{x}$ deviates manifold $\mathcal{M}$ and become the outlier, thereby demonstrates the reason for the incorrect decision made by $g$:
$g(f(\bm{x}+\bm{\delta}))=g(f(\bm{x})+e(\bm{\delta}))\neq g(f(\bm{x})),$ (6)
where $e(\bm{\delta})$ can be viewed as the deviation distance from the natural image manifold.
By this formulation, adversarial purification can be naturally seen as the process of estimating $e({\bm{\delta}})$ in the embedding space, the approximation of the natural image manifold. This means the adversarial embedding $f(\bm{x}_{\mathrm{adv}})=f(\bm{x}+\bm{\delta})$ can return the manifold along the same path $-e(\bm{\delta})$. While precise estimation of $e(\bm{\delta})$ is challenging, it is fortunate that only an approximation, denoted as $\tilde{e}(\bm{\delta})$ is sufficient to move the adversarial embedding back within the manifold.
Figure 2: (a) An illustration of activation of crucial semantic features on CIFAR-10 and ImageNet-1K. (b) Accuracy of different properties of images on three models, where the blurring operator is the median filter. Left: AutoAttack results on CIFAR-10. Middle and Right: AutoAttack on ImageNet.
3.2 Guided Shift
Considering adversarial purification as an inverse process to adversarial attacks, we can emulate Eq.(2) to formulate the optimization objective used to estimate $\tilde{e}(\bm{\delta})$:
$\min_{\bm{\delta}_{\mathrm{pfy}}}\ell(f\ \circ$ $g(\bm{x}_{\mathrm{adv}}+\bm{\delta}_{\mathrm{pfy}}),y)$ (7)
$\mathrm{s.t.}\ \ \ \|\bm{\delta}_{\mathrm{pfy}}\|\leq\epsilon_{\mathrm{pfy}}$ $\mathrm{and}\ e(\bm{\delta}_{\mathrm{pfy}})=-\tilde{e}(\bm{\delta})\approx-e(\bm{\delta}),$ (8)
where $\bm{\delta}_{\mathrm{pfy}}$ and $\epsilon_{\mathrm{pfy}}$ are defined to correspond to $\bm{\delta}$ and $\epsilon$ in Eq.(2) to offset the perturbation. However, even if $\epsilon_{\mathrm{adv}}$ can be considered as a hyperparameter, the ground-truth label $y$ is not accessible. External model-based adversarial purification [35, 51, 17, 30] typically train a purification model to minimize the global $\ell_{\mathrm{pfy}}$. Auxiliary function-based adversarial method [29, 38, 19] tend to design a suitable $\ell_{\mathrm{pfy}}$ to complete purification without relying on external generative models. They need to retrain the classifier or parameterized auxiliary function to effectively cooperate in removing adversarial perturbations, which means these methods all introduce parameters $\Theta$ in $\ell_{\mathrm{pfy}}$:
$\min_{\bm{\delta}_{\mathrm{pfy}}}\ell_{\mathrm{pfy}}(f(\bm{x}_{\mathrm{adv}}+\bm{\delta}_{\mathrm{pfy}});\Theta).\$ (9)
Based on our assumption in subsection 3.1 that adversarial images are outliers of the natural image manifold and the purification process can be considered as returning them to this manifold, we aim to design $\ell_{\mathrm{pfy}}$ without $\Theta$ by allowing them to shift towards the manifold adaptively:
$\min_{\bm{\delta}_{\mathrm{pfy}}}\ell^{\star}_{\mathrm{pfy}}(f(\bm{x}_{\mathrm{adv}}+\bm{\delta}_{\mathrm{pfy}}).$ (10)
We are starting to investigate whether a simple image transformation, such as color jitter, grayscale, Gaussian blur, solarization, and equalization can shift adversarial images towards the natural image manifold to improve classifier performance. We find that blurring adversarial examples can improve the robust accuracy of victim classifiers. Additionally, as shown in Fig 2(a), the blur operation can reactivate regions disrupted by adversarial images. This phenomenon suggests that a simple blurring operator can bring adversarial examples closer to the manifold and reverse adversarial properties.
Algorithm 1 Guided Shift
1:Adversarial example $\bm{x}_{\mathrm{adv}}$, iterations $T_{\mathrm{g}}$, step size $\eta_{1}$, a classifier $f$, purification bound $\epsilon_{\mathrm{pfy}}$, and the blurring operator $\mathrm{blur}(\cdot)$.
2:Guided result $\bm{x}_{\mathrm{g}}$
3:Random start $\bm{x}_{\mathrm{g}}^{0}\leftarrow\bm{x}_{\mathrm{adv}}+\varepsilon$
4:for $t=0,1,2,...,T_{\mathrm{g}}-1$ do
5:    $\bm{\tilde{x}}_{\mathrm{g}}^{t}\leftarrow\mathrm{blur}(\bm{x}_{\mathrm{g}}^{t})$
6:    $\bm{z}\leftarrow f(\bm{x}_{\mathrm{g}}^{t}),\bm{\tilde{z}}\leftarrow f(\bm{\tilde{x}}_{\mathrm{g}}^{t})$
7:    $\bm{x}_{\mathrm{g}}^{t+1}\leftarrow\bm{x}_{\mathrm{g}}^{t}+\eta_{1}\cdot\mathrm{sgn}\bigl{(}\nabla_{\bm{x}_{\mathrm{g}}^{t}}d(\bm{z},\bm{\tilde{z}})\bigr{)}$
8:    $\bm{x}_{\mathrm{g}}^{t+1}\leftarrow\mathrm{Clip}(\bm{x}^{t+1}_{\mathrm{g}},-\epsilon_{\mathrm{pfy}},\epsilon_{\mathrm{pfy}})$
9:    $\bm{x}_{\mathrm{g}}^{t+1}\leftarrow\mathrm{Clip}(\bm{x}^{t+1}_{\mathrm{g}},0,1)$
10:end for
11:$\bm{x}_{\mathrm{g}}\leftarrow\bm{x}_{\mathrm{g}}^{T_{\mathrm{g}}}$
Figure 3: Details of GS on CIFAR-10.
Therefore, we can accumulate this reversion to guide the adversarial examples back towards the manifold. For instance, we can pull the distance between adversarial examples and their blurred counterparts in the embedding space, which allows adversarial examples to converge closer to the natural image manifold. The distance of feature embeddings is defined by the Cosine Similarity:
$d(\bm{z}_{\mathrm{adv}},\bm{z}_{\mathrm{adv}}^{\prime})=\frac{\bm{z}_{\mathrm{adv}}\cdot\bm{z}_{\mathrm{adv}}^{\prime}}{\|\bm{z}_{\mathrm{adv}}\|\|\bm{z}_{\mathrm{adv}}^{\prime}\|},$ (11)
where $\bm{z}_{\mathrm{adv}}$ and $\bm{z}_{\mathrm{adv}}^{\prime}$ are the embeddings of adversarial examples and their blurred counterparts, respectively. we can shift the adversarial example $\bm{x}_{\mathrm{adv}}$ by the gradient of Eq.(11):
$\bm{x}_{\mathrm{adv}}^{+}=\bm{x}_{\mathrm{adv}}+\eta_{1}\cdot\nabla_{\bm{x}_{\mathrm{adv}}}d(f(\bm{x}_{\mathrm{adv}}),f(\bm{x}_{\mathrm{adv}}^{\prime})).$ (12)
Controlling the magnitude of the blurring applied carefully is essential to prevent excessive blurring that could render adversarial examples unrecognizable by the classifier. However, a small magnitude is not sufficient to shift images to reverse the distortion caused by adversarial perturbations. We therefore consider to move $\bm{x}_{\mathrm{adv}}$ by Eq.(12) with a small step size $\eta_{1}$ iteratively. This process is referred to as Guided Shift (GS). In each step of GS, we apply the same update rule:
$\bm{x}_{\mathrm{g}}^{t+1}=\bm{x}_{\mathrm{g}}^{t}+\eta_{1}$ $\cdot\nabla_{\bm{x}_{\mathrm{g}}^{t}}d(f(\bm{x}_{\mathrm{g}}^{t}),f(\bm{\tilde{x}}_{\mathrm{g}}^{t})),$ (13)
$\bm{\tilde{x}}_{\mathrm{g}}^{t+1}$ $=\mathrm{blur}(\bm{x}_{\mathrm{g}}^{t+1})$ (14)
where $\mathrm{blur}(\cdot)$ is the blurring operator and $\bm{x}_{\mathrm{g}}^{0}:=\bm{x}_{\mathrm{adv}}$, such as the median filter, Gaussian blur, and so on. The workflow of GS is shown in Algorithm 1, where we use $\mathrm{sgn}(\cdot)$ to regulate the step size.
To demonstrate the process of GS, we employ two blurring operators: the median filter and Gaussian blur to compute the cosine similarity between single-step result $\bm{x}_{\mathrm{g}}^{t}$ of GS and the natural image $\bm{x}$ ($\star$-$GS$-$N$), as well as between the blurred counterpart $\bm{\tilde{x}}_{\mathrm{g}}^{t}$ of $\bm{x}_{\mathrm{g}}^{t}$ and $\bm{x}$ ($\star$-$GS^{\prime}$-$N$). As shown in Fig. 3, in each step, the cosine similarity ($\star$-$GS$-$N$) is consistently smaller than the cosine similarity $\star$-$GS^{\prime}$-$N$. Furthermore, these cosine similarities consistently increase as GS progresses, and adversarial examples are gradually getting closer to the natural image manifold guided by their blurred counterparts.
3.3 Adaptive Projection
We also can observe an interesting phenomenon in Fig. 3. Regardless of the type of blurring operators, dash lines ($\star$-$GS^{\prime}$-$N$) and solid lines ($\star$-$GS$-$N$) tend to converge. Therefore, we can have the following assumption:
Assumption 1.
There is a limitation for Guided Shift in that the adversarial images can be guided by their blurred counterparts only toward the manifold rather than fully back to the manifold. In other words, this indicates that Guided Shift approximates a convex function.
(a) AP w/o. RP in Adv $\bm{x}_{\mathrm{init}}$
(b) AP in Adv $\bm{x}_{\mathrm{init}}$
(c) AP w/o. RP in Nat $\bm{x}_{\mathrm{init}}$
(d) AP in Nat $\bm{x}_{\mathrm{init}}$
Figure 4: Intutive explanation of AP in the context of adversarial or natural $\bm{x}_{\mathrm{init}}$. PR: the perceptual regularization (Eq.(20)).
Algorithm 2 Adaptive Projection
1:Input image $\bm{x}_{\mathrm{init}}$, guided result $\bm{x}_{\mathrm{g}}$, iterations $T_{\mathrm{p}}$, a classifier $f$, purification bound $\epsilon_{\mathrm{pfy}}$ , and candidate layers $S$.
2:Adaptive result $\bm{x}_{\mathrm{p}}$
3:Initialize $\bm{x}_{\mathrm{p}}^{0}\leftarrow\bm{x}_{\mathrm{init}}$
4:Set step size $\eta_{2}\leftarrow\epsilon_{\mathrm{pfy}}/T_{\mathrm{p}}$
5:for $t=0,1,2,...,T_{\mathrm{p}}-1$ do
6:    for $l\in S$ do
7:        $\Delta\bm{u}_{\mathrm{g}}^{l}\leftarrow f_{l}(\bm{x}_{\mathrm{g}})-f_{l}(\bm{x}_{\mathrm{init}})$
8:        $\Delta\bm{u}_{\mathrm{p}}^{l}\leftarrow f_{l}(\bm{x}_{\mathrm{p}}^{t})-f_{l}(\bm{x}_{\mathrm{init}})$
9:        $\mathcal{L}\leftarrow\mathcal{L}+(-\Delta\bm{u}_{\mathrm{g}}^{l}\cdot\Delta\bm{u}_{\mathrm{p}}^{l})$
10:    end for
11:    $\mathcal{L}\leftarrow\frac{\lambda_{1}}{\|S\|}\mathcal{L}+\lambda_{2}\|\phi(\bm{x}_{\mathrm{p}}^{t}-\phi(\bm{x}_{\mathrm{init}}))\|_{2}$
12:    $\bm{x}_{\mathrm{p}}^{t+1}\leftarrow\bm{x}_{\mathrm{p}}^{t}-\eta_{2}\cdot\mathrm{sgn}(\nabla_{\bm{x}_{\mathrm{p}}^{t}}\mathcal{L})$
13:    $\bm{x}_{\mathrm{p}}^{t+1}\leftarrow\mathrm{Clip}(\bm{x}_{\mathrm{p}}^{t+1},-\epsilon_{\mathrm{pfy}},\epsilon_{\mathrm{pfy}})$
14:    $\bm{x}_{\mathrm{p}}^{t+1}\leftarrow\mathrm{Clip}(\bm{x}_{\mathrm{p}}^{t+1},0,1)$
15:end for
16:$\bm{x}_{\mathrm{p}}\leftarrow\bm{x}_{\mathrm{p}}^{T_{\mathrm{p}}}$
Certainly, this limitation arises from the exceedingly blurred images whose embeddings cannot be recognized correctly by the classifier. To fully return adversarial images to the manifold, we must overcome this limitation. A natural intuition is to allow current images to move adaptively instead of being guided by blurred counterparts, which requires us to target a basic direction for current images. Fortunately, the result of Guided Shift already provides this basic direction. Therefore, we propose Adaptive Projection (AP) that can be defined as:
$\max_{\bm{x}_{\mathrm{p}}}\lambda_{1}\frac{1}{\|S\|}\sum_{l\in S}\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})-\lambda_{2}\|\phi(\bm{x}_{\mathrm{p}})-\phi(\bm{x}_{\mathrm{init}})\|_{2}\ \ \ \ \mathrm{s.t.}\ \ \ \|\bm{x}_{\mathrm{p}}-\bm{x}_{\mathrm{init}}\|\leq\bm{\delta}_{\mathrm{pfy}},$ (15)
where $S\subseteq L$ as a candidate in the set $L=\{l_{1},l_{2},...,l_{m}\}$ of a $m$ layers model $f$, $\|S\|$ denotes the number of elements of the set $S$. $\bm{x}_{\mathrm{init}}$ represents the input image (i.e., $\bm{x}_{\mathrm{init}}=\bm{x}_{\mathrm{adv}}$ when the input image is adversarial). For the $l^{th}$ layer of classifiers, the first term of Eq. (15) can be written as:
$\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})=-\Delta\bm{u}^{l}_{\mathrm{p}}\cdot\Delta\bm{u}^{l}_{\mathrm{g}},$ (16)
where $\bm{x}_{\mathrm{p}}$ is the final result, $\Delta\bm{u}^{l}_{\mathrm{p}}$ and $\Delta\bm{u}^{l}_{\mathrm{g}}$ are two vectors of flattened feature maps defined as follow:
$\Delta\bm{u}^{l}_{\mathrm{p}}=f_{l}(\bm{x}_{\mathrm{p}})-f_{l}(\bm{x}_{\mathrm{init}}),\ \ \ \Delta\bm{u}^{l}_{\mathrm{g}}=f_{l}(\bm{x}_{\mathrm{g}})-f_{l}(\bm{x}_{\mathrm{init}}),$ (17)
where $f_{l}$ denoted as feature maps at layer $l$ of the classifier, $\bm{x}_{\mathrm{p}}$ is initialized by $\bm{x}_{\mathrm{init}}$. Maximizing Eq.(16) is equivalent to maximizing the projection of $\bm{u}^{l}_{\mathrm{p}}$ onto $\bm{u}^{l}_{\mathrm{g}}$ since $\|\bm{u}^{l}_{\mathrm{g}}\|$ is a constant. The increase in projection implies that $\bm{x}_{\mathrm{p}}$ is not restricted by the blurred images to continue moving along the direction of the result of GS, which allows $\bm{x}_{\mathrm{p}}$ move independently toward the natural image manifold.
However, We typically do not know whether an image is adversarial in real-world applications (i.e., $\bm{x}_{\mathrm{init}}=\bm{x}$ when $\bm{x}$ is natural), repeatedly guiding images by blurred counterparts may cause natural images to deviate from the manifold. As shown in Fig. 4, this deviation will be enhanced by AP. Therefore, we introduce a perceptual regularization term in Eq.(15). Let $F_{1}$ is the dynamic of Eq.(16):
$F_{1}=\nabla_{\bm{x}_{\mathrm{p}}}\frac{1}{\|S\|}\sum_{l\in S}\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})\ ,$ (18)
If there exists a momentum $F_{2}$, such that:
$\left\{\begin{array}[]{lll}F_{1}>F_{2},&\bm{x}_{\mathrm{init}}\ is\ adversarial,\\ F_{1}<F_{2}&\bm{x}_{\mathrm{init}}\ is\ natural,\\ \end{array}\right.$ (19)
then maximizing Eq. (16) allows the result $\bm{x}_{p}$ to locate in the natural image manifold. In this work, the dynamic $F_{2}$ can be computed by LPIPS distance [55]. The LPIPS distance $d(\bm{x}_{1},\bm{x}_{2})$ between images $\bm{x}_{1}$ and $\bm{x}_{2}$ is then defined as $d(\bm{x}_{1},\bm{x}_{2})\triangleq\|\phi(\bm{x}_{1})-\phi(\bm{x}_{2})\|_{2}$. Let $\hat{f}(\bm{x})$ denote channel-normalized activations at the $l$-th layer of the classifier. Then, $\phi(\bm{x})\triangleq(\frac{\hat{f}(\bm{x})}{\sqrt{w_{1}h_{1}}},...,\frac{\hat{f}(\bm{x})}{\sqrt{w_{m}h_{m}}})$, where $w_{l}$ and $h_{l}$ are the width and height of activations of layer $l$, respectively.
Finally, the momentum $F_{2}$ can be written as:
$\centering F_{2}=-\lambda_{2}\cdot\nabla_{\bm{x}_{\mathrm{p}}}\|\phi(\bm{x}_{\mathrm{p}})-\phi(\bm{x}_{\mathrm{init}})\|_{2}.\@add@centering$ (20)
In contrast to the original LPIPS implementation, we do not require this distance to approximate human perceptual judgments. Instead, we aim for the perceptual distance between $\bm{x}_{\mathrm{g}}$ and $\bm{x}_{\mathrm{init}}$ based on the classifier itself to be small. Hence, we can directly use $d(\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{init}})$ to evaluate the distance from the perspective of the classifier. The workflow of AP is shown in Algorithm 2.
4 Experiments
Datasets and Metrics. Three benchmarks CIFAR-10 [22], CIFAR-100, and ImageNet-1K [9] are considered to evaluate our method. We compare our method with the state-of-the-art adversarial training methods reported in standard benchmark RobustBench [7] and other adversarial purification methods [38, 29, 19, 41, 16, 51, 43, 30, 17, 26, 45]. In all experiments, we consider two metrics to evaluate the performance of all methods: standard accuracy and robust accuracy. The standard accuracy measures the performance of the defense method on natural images, while the robust accuracy measures the performance on adversarial images.
Adversarial attacks and victim classifiers. We evaluate our methods with three attacks: AutoAttack [6], DI2-FGSM [49], and BPDA [3]. AutoAttack is a powerful adaptive attack commonly used in most defense studies. We also use DI2-FGSM to seek whether an attack robust to the blurring operator can affect our method. Additionally, we consider BPDA to attack our purification module in the worst-case scenario. These attacks will attack three victim classifiers including ResNet-18 [14], ResNet-50, and WideResNet-28-10 [53].
Table 1: Standard and robust accuracy (%) against AutoAttack $\ell_{\infty}(\epsilon=8/255)$ in comparison with AT &\&& ABP methods. The first section corresponds to AT methods and the second to ABP methods. $\dagger$ Since Hwang et al. [19] only reports the performance of WideResNet-34-10 in the original paper, and the pre-trained parameterized auxiliary checkpoint has been removed, we use this result for comparison.
Require Training ResNet-18 WideResNet-28-10
Classifier Auxiliary Method Standard Robust Method Standard Robust
CIFAR-10
(Gowal et al., 2021) 87.35 59.12 (Gowal et al., 2021) 87.50 63.99
(Sehwag et al., 2021) 84.59 56.19 (Pang et al., 2022) 88.61 61.40
(Rade et al., 2021) 89.02 58.17 (Xu et al., 2023) 93.69 65.62
(Addepalli et al., 2022) 85.71 52.90 (Wang et al., 2023) 92.44 67.31
[…]
(Mao et al., 2021) - 58.20 (Mao et al., 2021) - 67.15
(Hwang et al., 2023$\dagger$ 87.02 56.63 (Hwang et al., 2023) 87.02 56.63
GS 49.56 58.82 GS 50.24 59.68
GS+AP (ZeroPur) 92.56 69.62 GS+AP (ZeroPur) 91.81 68.60
CIFAR-100
(Rade et al., 2021) 61.50 29.50 (Pang et al., 2022) 63.66 31.67
(Addepalli et al., 2022) 65.45 28.58 (Rebuffi et al., 2021) 62.41 33.03
[…]
(Mao et al., 2021) - 25.45 (Mao et al., 2021) - 33.16
(Hwang et al., 2023) 64.73 32.86 (Hwang et al., 2023) 64.73 32.86
GS 30.42 34.35 GS 26.56 32.06
GS+AP (ZeroPur) 52.45 41.42 GS+AP (ZeroPur) 64.88 34.58
4.1 Quantitative Evaluation and Comparison
For adversarial defense, adversarial training (AT) involves adversarial examples in the classifier training, and auxiliary-based purification (ABP) introduces an auxiliary function to cooperate with the victim classifier. These methods do not rely on external generative models but still necessitate training auxiliary functions or retraining classifiers. External model-based purification (EBP) fine-tunes or retrains an external generative model to remove adversarial perturbations in adversarial images. Since the performance of EBP is ensured by the modeling capability of generative models, it typically outperforms ABP. In this section, for a fair comparison, we will first compare our method with both AT and ABP and then compare it with EBP. ’Require Training’ in Table 1,  3 and  3 denotes that the methods require retraining external generative models or auxiliary functions or victim classifiers.
Comparison with AT &\&& ABP. Table 1 reports the standard and robust accuracy against AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$). ZeroPur achieves an approximate $10\%$ increase in robust accuracy for ResNet-18, and this improvement surpasses the performance of most other methods using WideResNet-28-10. We also report the comparison on ImageNet1-K in Table 3, where ZeroPur improves robust accuracy by $30\%$ compared to AT and ABP methods. These results also demonstrate that AP can significantly move adversarial images toward the natural image manifold. It is noteworthy to mention that our method consistently presents better performance in small architectures such as ResNet-18 compared to large architectures. This stems from the inclination of our method to ‘fine-tune’ adversarial images, enabling them to approach natural images, whereas larger models often introduce more hindrances to this process.
Comparison with EBP. Training an external generative model to model a transformation from adversarial images to natural images can contribute to an outstanding purification method, but this process is exceedingly computation-consuming. Table 3 and Table 3 report the comparison result between ZeroPur and EBP. The robust accuracy achieved by ZeroPur is only slightly lower than DISCO (Ho et al., 2022) which involves training an external model on ImageNet. Therefore, adjusting ZeroPur a little bit (see the first part of Subsection 4.2) can achieve comparable performance with these EBP methods. Other detailed results are reported in Appendix A.4.
Table 2: Standard and robust accuracy (%) against AutoAttack $\ell_{\infty}(\epsilon=8/255)$ on CIFAR-10 in comparison with EBP methods.
Require Training WideResNet-28-10
Cls Ext Method Standard Robust
(Sun et al., 2019) 82.22 67.92
(Hill et al., 2020) 84.12 78.91
(Yoon et al., 2021) 86.14 80.24
(Ughini et al., 2022) - 59.57
(Nie et al., 2022) 89.02 70.64
(Ho et al., 2022) 89.26 85.56
(Lin et al., 2024) 90.62 72.85
[…]
ZeroPur 85.02 82.76
Table 3: Standard and robust accuracy (%) against AutoAttack $\ell_{\infty}(\epsilon=4/255)$ on ImageNet-1K.
Training ResNet-50
Cls Aux Ext Method Standard Robust
Comparison with AT &\&& ABP
(Salman et al., 2020) 64.02 34.96
(Wong et al., 2020) 55.62 26.24
(Bai et al., 2021) 67.38 35.51
[…]
ZeroPur 61.05 63.09
Comparison with EBP
(Wang et al., 2022) 70.17 68.782
(Nie et al., 2022) 67.79 40.93
(Ho et al., 2022) 71.22 69.521
[…]
4.2 Discussion on ZeroPur
Evaluating ZeroPur on protecting the victim classifier trained with different data augmentation. To further discuss the performance of ZeroPur, we consider three data augmentation strategies (’Vanilla’, ‘Base’, ‘Strong’) to train victim classifiers (See Appendix A.2 for details). ‘Base’ is the most common data augmentation in training classifiers. In the experiment, ZeroPur using a median filter ($3\times 3$ window size) is applied in ’Vanilla’ classifiers and Gaussian blur ($\sigma=1.2$) in ‘Base’ and ’Strong’ classifiers. Table 5 reports the standard and robust accuracy of ZeroPur. We can see that ’Strong’ data augmentation benefits our method. The reason is that such augmentations enable the blurred counterparts to guide adversarial images toward the natural image manifold more effectively, which contributes to precise adaptive projection. The performance of ZeroPur in Table 3 also demonstrates this phenomenon.
[…]
Accuracy ResNet-18 WRN-28-10
Vanilla Base Strong Vanilla Base Strong
CIFAR-10
Standard 82.04 92.56 90.05 91.22 91.81 85.02
Robust 59.38 69.62 83.84 59.16 68.60 82.76
CIFAR-100
Standard 55.45 52.45 67.11 64.89 64.88 66.39
Robust 30.42 41.42 50.66 30.81 34.58 50.25
Table 5: The ablation study on CIFAR-10 using perceptual regularization.
Impact of perceptual regularization. The important of incorporating the perceptual regularization in (Eq.(15)) is revealed by Fig. 5. We can see the robust accuracy of AP experiences a slight decrease in (a) and (b), while the standard accuracy shows significant improvement in (c) and (d). This phenomenon is observed in both ResNet-18 and WideResNet-28-10, indicating the effectiveness of perceptual regularization in AP.
ZeroPur defense against adversarial attacks robust to blurring. ZeroPur utilizes the blurring operator to return adversarial images to the natural image manifold. However, certain attacks, such as DI2-FGSM [49], inherently withstand blurring operations. Indeed, they cannot be entirely immune to ZeroPur. As shown in Table 7, replacing the blurring operator with TVM [13] leads to improved purification. The robust accuracy against AutoAttack increases to $75.96\%$ without training, which outperforms the optimal performance of GS(Blur)+AP by $6.34\%$. This result also demonstrates that our method can enhance any operator capable of destroying adversarial properties, not just blurring.
ZeroPur defense against attacks that bypass the purification module. The efficiency of purification methods can be challenged by the BPDA attack that approximates gradients, bypassing the purification module. Table 7 reports the performance of ZeroPur defense against BPDA attack, where the victim classifiers are WideResNet-28-10 (WRN) and ResNet-18 (R18). We can see that the robust accuracy of ZeroPur decreases, but it remains stable, which suggests that ZeroPur still offers acceptable defense. Since the natural accuracy of ’Vanilla’ WRN on natural images is $71.50\%$, while that of ’Vanilla’ R18 is $57.01\%$, which may result in an approximately $20\%$ gap in evaluation on BPDA (See Table 8 in Appendix A.2).
[…]
Method DI2-FGSM AutoAttack
Vanilla Base Strong Vanilla Base Strong
GS (Blur) 35.77 40.96 66.66 54.37 58.82 79.36
GS (TVM) 38.08 46.79 45.93 56.25 67.81 68.48
GS (Blur) + AP 35.34 59.81 84.41 59.38 69.62 83.84
GS (TVM) + AP 50.51 65.98 66.74 61.75 75.96 78.98
[…]
Model PGD-10 PGD-20 PGD-40
Vanilla WRN 50.61 50.71 50.54
R18 34.44 34.88 32.18
Base WRN 34.16 33.38 34.42
R18 39.88 39.90 39.48
Strong WRN 68.84 69.45 68.50
R18 70.65 70.45 70.43
5 Conclusion
We propose a succinct training-free method for adversarial purification, named ZeroPur. Our method significantly outperforms previous state-of-the-art adversarial training and auxiliary-based purification methods, while demonstrating comparability with external model-based purification methods. Despite the improvements, ZeroPur has a major limitation: the absence of external models restricts our ability to enhance defense against strong adaptive attacks. To overcome this challenge, it is necessary to design a module that emulates recent works on purification by diffusion models capable of effectively defending BPDA.
Appendix A Appendix
A.1 Proofs
Theorem 1.
Suppose $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ is twice differentiable at point $\bm{x}$. Let $e_{1}=\nabla f_{1}(\bm{x})^{T}\bm{\delta}+\frac{1}{2}\bm{\delta}^{T}\nabla^{2}f_{1}(\bar{\bm{x}})\bm{\delta}$, and there exists $w_{1}\in[0,1]$ such that $\bar{\bm{x}}=w_{1}\bm{x}+(1-w_{1})(\bm{x}+\bm{\delta})$, then we have the forward of $f$:
$f_{1...l}(\bm{x}+\bm{\delta})$ $=f_{1...l}(\bm{x})+e_{l}(\bm{\delta}),$ (1)
$where\ \ \ e_{l}(\bm{\delta})=\nabla f_{l}(f_{1...l-1}(\bm{x}))^{T}$ $e_{l-1}(\bm{\delta})+\frac{1}{2}e_{l-1}(\bm{\delta})^{T}\nabla^{2}\bar{f}_{1...l-1}(\bm{x})e_{l-1}(\bm{\delta}).$ (2)
where there exists $w_{l}\in[0,1]$ such that $\bar{f}_{1...l-1}(\bm{x})=w_{l}f_{1...l-1}(\bm{x})+(1-w_{l})(f_{1...l-1}(\bm{x})+e_{l-1}(\bm{\delta}))$.
Proof.
By a first-order Taylor series, we have:
$f(\bm{x})=f(\bm{x}_{0})+\nabla f(\bm{x}_{0})^{T}(\bm{x}-\bm{x}_{0})+\frac{1}{2}(\bm{x}-\bm{x}_{0})^{T}\nabla^{2}f(\bar{\bm{x}})(\bm{x}-\bm{x}_{0}).$ (3)
There exists $w\in[0,1]$ such that $\bar{\bm{x}}=w\bm{x}+(1-w)\bm{x}_{0}$. for the forward of an adversarial example $\bm{x}+\bm{\delta}$ in the first layer $f_{1}$, we have:
$f_{1}(\bm{x}+\bm{\delta})=f_{1}(\bm{x})+\underbrace{{\nabla f_{1}(\bm{x})^{T}\bm{\delta}+\frac{1}{2}\delta^{T}\nabla^{2}f_{1}(\bar{\bm{x}})\bm{\delta}}}_{e_{1}(\bm{\delta})}.$ (4)
Plugging $f_{2}$ into Eq.(4), we also have:
$f_{1,2}(\bm{x}+\bm{\delta})$ $=f_{2}(f_{1}(\bm{x}+\bm{\delta}))=f_{2}(f_{1}(\bm{x})+e_{1}(\bm{\delta}))$ (5)
$=\underbrace{f_{2}(f_{1}(\bm{x}))}_{f_{1,2}(\bm{x})}+\underbrace{\nabla f_{2}(f_{1}(\bm{x}))^{T}e_{1}(\bm{\delta})+\frac{1}{2}e_{1}(\bm{\delta})^{T}\nabla^{2}\bar{f}_{1}(\bm{x})e_{1}(\delta)}_{e_{2}(\bm{\delta})}.$ (6)
when $\bm{x}+\bm{\delta}$ is forward through the $l$-th layer, we have Theorem 1. $\square$
A.2 Implementation Details of Classifiers
Our experiments on CIFAR-10 and CIFAR-100 included ResNet-18 and WideResNet-28-10 trained on three different data augmentation strategies ‘Vanilla’, ‘Base’, and ‘Strong’. Table 8 shows the training settings and natural accuracy. ‘Base’ is the most common data augmentation used by adversarial training methods and adversarial purification methods for better natural accuracy.
Victim Classifier Case Trained On Natural Accuracy (%)
ReCrop. ColorJ. GrayS. GauBlur. Solar. Equal. HorFlip. CIFAR-10 CIFAR-100
ResNet-18 Vanilla 83.80 57.01
Base 93.10 71.58
Strong 91.08 68.52
WRN-28-10 Vanilla 91.34 71.50
Base 93.83 74.95
Strong 91.09 67.52
Table 8: The adopted data augmentation and natural accuracy (%) of victim classifiers. The enumerated data augmentation are, in order, ResizeCrop, ColorJitter, Grayscale, Solarization, Equalization, and HorizontalFlip.
All classifiers were trained with the SGD optimizer with a cosine decay learning rate schedule and a linear warm-up period of 10 epochs. The weight decay is $5.0\times 10^{-4}$ and the momentum is $0.9$. The initial learning rate is set to $0.1$. Classifiers were trained for 120 epochs on 4 Tesla V100 GPUs, where the batch size is 512 per GPU for ResNet-18 and 128 per GPU for WideResNet-28-10.
A.3 Implementation Details of Adversarial Attacks
AutoAttack [6].  We use AutoAttack to compare with the start-of-the-art methods. The robust classifier for adversarial training methods provided by RobustBench [7] benchmark available at https://robustbench.github.io. The code for adversarial purification methods is provided by their respective papers.
There are two versions of AutoAttack: (i) the Standard including AGPD-CE, AGPD-T, FAB-T, and Square, and (ii) the Rand version including APGD-CE and APGD-DLR. Considering that most of the adversarial purifications choose the Rand version, all the performance in this work we report is also in the Rand version. Code is available at https://github.com/fra31/auto-attack.
DI2-FGSM [49]. DI2-FGSM crafts adversarial examples by applying various transformations, enhancing their robustness against blurring operations. We use this attack to evaluate ZeroPur in Section 4.2, implemented by torchattacks [21]. Code is available at https://github.com/Harry24k/adversarial-attacks-pytorch.
BPDA [3]. We use BPDA, approximating the gradient of the purifier module (ZeroPur) as 1 during the backward pass. The 10, 20, and 40 iterations are applied in our experiment. Other settings are the same as those used in PGD Attack.
A.4 More Experimental Results
In Table 9 and Table 10, we compare our method with AT, ABP, and EBP methods on CIFAR-10 against $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$ threat model, detailing training requirements for each method. We also report the performance of ZeroPur when victim classifiers use strong data augmentation (ZeroPur). Comparing all methods, ZeroPur always achieves the second-best performance. However, when comparing ABP methods, ZeroPur achieves optimal performance.
Table 9: Standard and robust accuracy (%) against AutoAttack $\ell_{\infty}(\epsilon=8/255)$ on CIFAR-10 in comparison with three types methods, obtained by different classifier architectures.
Defense Type Method Training Architecture Standard Acc Robust Acc
AT (Gowal et al., 2021) Classifier ResNet-18 87.35 59.12
(Gowal et al., 2021) Classifier WRN-28-10 87.50 63.99
(Sehwag et al., 2021) Classifier ResNet-18 84.59 56.19
(Rade et al., 2021) Classifier ResNet-18 89.02 58.17
(Addepalli et al., 2022) Classifier ResNet-18 85.71 82.90
(Pang et al., 2022) Classifier WRN-28-10 88.61 61.40
(Xu et al., 2023) Classifier WRN-28-10 93.69 65.62
(Wang et al., 2023) Classifier WRN-28-10 92.44 67.31
[…]
(Hill et al., 2020) EBM+LD WRN-28-10 84.12 78.91
(Yoon et al., 2021) DSM+LD WRN-28-10 86.14 80.24
(Ughini et al., 2022) DeepFill [52] WRN-28-10 - 59.57
(Nie et al., 2022) DDPM [18] WRN-28-10 89.02 70.64
(Nie et al., 2022) DDPM [18] WRN-70-16 90.07 71.29
(Ho et al., 2022) LIIF [5] WRN-28-10 89.26 85.56
(Lin et al., 2024) MAE [15] WRN-28-10 90.62 72.85
(Lin et al., 2024) MAE [15] WRN-70-16 91.99 76.37
[…]
(Shi et al., 2021) Classifier WRN-28-10 91.89 68.56
(Mao et al., 2021) Auxiliary branch ResNet-18 - 58.20
(Mao et al., 2021) Auxiliary branch WRN-28-10 - 67.15
(Hwang et al., 2023) Auxiliary branch WRN-34-10 87.02 56.63
ZeroPur N/A ResNet-18 92.56 69.62
ZeroPur N/A ResNet-18 90.05 83.84
ZeroPur N/A WRN-28-10 91.81 68.60
ZeroPur N/A WRN-28-10 85.02 82.76
Table 10: Standard and robust accuracy (%) against AutoAttack $\ell_{2}(\epsilon=0.5)$ on CIFAR-10 in comparison with three types methods, obtained by different classifier architectures.
Defense Type Method Training Architecture Standard Acc Robust Acc
AT (Rebuffi et al., 2021) Classifier ResNet-18 90.33 75.86
(Sehwag et al., 2021) Classifier ResNet-18 89.76 74.41
(Wang et al., 2023) Classifier WRN-28-10 95.16 83.68
(Rebuffi et al., 2021) Classifier WRN-28-10 91.79 78.80
(Sehwag et al., 2021) Classifier WRN-34-10 90.93 77.24
[…]
(Ughini et al., 2022) DeepFill [52] WRN-28-10 - 45.12
(Nie et al., 2022) DDPM [18] WRN-28-10 91.03 78.58
(Nie et al., 2022) DDPM [18] WRN-70-16 92.68 80.60
(Ho et al., 2022) LIIF [5] WRN-28-10 89.26 88.47
(Lin et al., 2024) MAE [15] WRN-28-10 90.62 80.47
(Lin et al., 2024) MAE [15] WRN-70-16 91.99 81.35
[…]
ZeroPur N/A ResNet-18 90.77 86.56
ZeroPur N/A WRN-28-10 72.38 72.12
ZeroPur N/A WRN-28-10 81.52 86.61
[…]
Operator Level ResNet-18 WRN-28-10
Vanilla Base Strong Vanilla Base Strong
Median $3\times 3$ 59.38 72.27 60.74 59.16 67.56 70.01
$5\times 5$ 34.17 64.81 44.63 39.81 58.41 50.49
$7\times 7$ 27.25 54.41 34.68 28.94 48.26 40.58
Gaussian $\sigma=0.6$ 38.82 65.43 64.92 53.21 58.31 73.79
$\sigma=1.2$ 26.20 69.62 83.84 27.97 68.60 82.76
$\sigma=1.8$ 20.45 60.67 67.89 26.82 59.91 78.82
[…]
Operator Level ResNet-18 WRN-28-10
Vanilla Base Strong Vanilla Base Strong
Median $3\times 3$ 82.04 92.34 90.90 91.22 91.81 90.65
$5\times 5$ 83.73 91.87 90.92 91.06 91.34 88.97
$7\times 7$ 83.66 91.56 90.87 90.97 90.34 87.70
Gaussian $\sigma=0.6$ 83.57 91.64 90.72 91.09 90.02 90.68
$\sigma=1.2$ 83.63 92.56 90.05 91.14 91.81 85.02
$\sigma=1.8$ 83.69 92.36 90.79 51.26 35.95 80.27
Table 11 and Table 12 report the robust accuracy and standard accuracy of ZeroPur using different blur operators in GS. We use 5 different blurring operators including median filters with windows size $3\times 3$, $5\times 5$, and $7\times 7$ and Gaussian blurring kernel with $\sigma=0.6$, $1.2$, and $1.8$. The results show that if victim classifiers are not trained on strong data augmentation, low-quality embeddings of images from excessive blurring will hinder ZeroPur’s ability to move adversarial images. On the contrary, the victim classifiers trained on strong data augmentation can recognize these embeddings, enabling ZeroPur to provide precise movement direction.
We provide visual examples in Fig. 5 to illustrate several results and the visualization of this process, where diff_$I_{a}$_$I_{b}$ = $\mathrm{abs}(I_{a}-I_{b})$ represents the difference between the two images. The resemblance between diff_gs_nat and diff_adv_nat outlines suggests that Guided Shift can effectively reverse the adversarial perturbation at a coarse level. Furthermore, the close correspondence in detail between diff_adv_nat and diff_ap_nat demonstrates the precise approximation of this perturbation by Adaptive Projection. diff_ap_adv quantifies the distance of the adversarial image moved by ZeroPur.
(a) Examples purified by GS
(b) Examples purified by GS + AP
Figure 5: Visual examples of ZeroPur against $\ell_{\infty}$ threat model ($\epsilon=4/255$) on ImageNet-1K. The central two columns depict the results of GS and AP, while the red label denotes error prediction and the green label denotes correct prediction. diff_$I_{a}$_$I_{b}$ = $\mathrm{abs}(I_{a}-I_{b})$ represents the difference between the two images.