### v2

# ZeroPur: Succinct Training-Free Adversarial Purification

Abstract
Adversarial purification is a kind of defense technique that can defend against various unseen adversarial attacks without modifying the victim classifier. Existing methods often depend on external generative models or cooperation between auxiliary functions and victim classifiers. However, retraining generative models, auxiliary functions, or victim classifiers relies on the domain of the fine-tuned dataset and is computation-consuming. In this work, we suppose that adversarial images are outliers of the natural image manifold, and the purification process can be considered as returning them to this manifold. Following this assumption, we present a simple adversarial purification method without further training to purify adversarial images, called ZeroPur. ZeroPur contains two steps: given an adversarial example, Guided Shift obtains the shifted embedding of the adversarial example by the guidance of its blurred counterparts; after that, Adaptive Projection constructs a directional vector by this shifted embedding to provide momentum, projecting adversarial images onto the manifold adaptively. ZeroPur is independent of external models and requires no retraining of victim classifiers or auxiliary functions, relying solely on victim classifiers themselves to achieve purification. Extensive experiments on three datasets (CIFAR-10, CIFAR-100, and ImageNet-1K) using various classifier architectures (ResNet, WideResNet) demonstrate that our method achieves state-of-the-art robust performance. The code will be publicly available.
1 Introduction
Over the past decade, a substantial body of work has confirmed that deep neural networks (DNNs) are highly vulnerable to imperceptible adversarial perturbations that cause wrong decisions [47, 10, 32, 4, 54]. This potential vulnerability, despite their remarkable performance, poses a significant challenge for security-critical applications. Thus, designing effective and efficient adversarial defense techniques is essential to ensure the reliability and trustworthiness of DNNs in real-world scenarios.
One kind of adversarial defense technique is adversarial training [30, 59, 28, 23], which involves adversarial examples in the model training, enabling the model to adapt adversarial perturbations empirically. However, these approaches typically require huge computational resources [41, 53] and suffer from performance degradation [24] in the presence of unseen attacks that are not involved in training. This limitation hinders the application of adversarial training in realistic scenarios.
Figure 1: An illustration of ZeroPur.
Different from adversarial training, adversarial purification [56, 42, 33, 31] aims to remove adversarial perturbations in adversarial examples. These methods do not require adversarial examples in the model training and effectively defend against unseen attacks, making them more applicable in real-world scenarios. However, the existing purification methods often depend on external generative models [9, 45] or cooperation between external auxiliary functions and the victim classifier [42, 31, 17]. Retraining generative models, parameterized auxiliary functions, or the classifier relying on the domain of the fine-tuned dataset is computationally demanding and restricts their flexibility.
Inspired by the natural image manifold hypothesis, we suppose that adversarial images are outliers of the natural image manifold, and the purification process aims to return them to this manifold. We propose a simple adversarial purification method called ZeroPur, where ’Zero’ means our method does not need to train any external models, parameterized auxiliary functions, and victim classifiers. As illustrated in Figure 1, ZeroPur comprises two stages: Guided Shift (GS) and Adaptive Projection (AP) for purifying the adversarial image. Specifically, we observe that the blurring operator provides a reasonable direction towards natural image on the manifold for an adversarial sample. Therefore, GS iteratively guides adversarial sample towards its blurred counterparts, obtaining shifted embedding that approximate the manifold. Due to the inherent limitations of low-quality embeddings from blurred counterparts, adversarial sample may not precisely return to the target location, but the shifted embedding after multiple iterations provide a more reliable direction. Subsequently, we introduce AP to construct a directional vector based on this shifted embedding and project the adversarial image adaptively onto the manifold by maximizing the projection along this direction. Despite its embarrassingly simple design, ZeroPur consistently outperforms most state-of-the-art (SOTA) adversarial training and purification methods.
The main contributions of the current work are as follows:
  • We analyze the relationship between adversarial attack and adversarial purification based on the natural image manifold hypothesis, and show that a simple blurring operator can bring adversarial examples that are out of the natural image manifold closer to the manifold.
  • We present a succinct adversarial purification approach named ZeroPur, including two stages: Guided Shift and Adaptive Projection, which requires no retraining of external models, parameterized functions, and the victim classifier.
  • Extensive experiments demonstrate that the proposed approach outperforms most state-of-the-art auxiliary-based adversarial purification methods and achieves competitive performance compared to other external model-based purification methods.
2 Related Work
Adversarial training (AT). AT [30, 59, 28, 23] improves the robustness of DNNs by integrating adversarial examples into the training data and reformulating the optimization objective. However, the computational cost of adversarial training is significantly huge due to the necessity of repeatedly performing backpropagation to craft adversarial examples. While recent works investigate reducing the time cost of adversarial training, they are still restricted by issues such as low robust performance [41, 37] and other unexpected results [52, 50] (i.e., catastrophic overfitting). Moreover, even if a model is robust against a specific set of known attacks, it is still fragile against other unseen attacks that were not involved in training [19, 24, 7].
External model-based adversarial purification (EBP). Samangouei et al. [39] propose defense-GAN, a generator that models the distribution of natural images, which enables the transformation from adversarial examples to natural images. Song et al. [44] assume that adversarial examples primarily reside in the low probability density region of the training distribution, and design PixelDefend to approximate this distribution using the PixelCNN [49]. Later, utilizing score-based models [56] and diffusion models [43, 33, 25, 18] as purification models is introduced and shown to achieve significantly improved robust performance. Recently, Lin et al. [27] propose a framework called AToP that fine-tunes the purification model adversarially to integrate the benefits of both adversarial training and adversarial purification. However, these works rely on external generative models, substantial datasets (e.g., crafting adversarial examples), and computational resources.
Auxiliary-based adversarial purification (ABP). In contrast to the above two approaches, ABP tends to introduce an auxiliary function to cooperate with the classifier to purify adversarial images. Shi et al. [42] propose a lightweight purification method, SOAP, which uses self-supervised loss to realize online purification. SOAP no longer depends on generative models but requires classifiers to incorporate the corresponding auxiliary loss (self-supervised loss) in the training stage. To further reduce the time cost of purification, recent works introduce parameterized auxiliary functions, allowing purification to be accomplished through the exclusive training of these functions. Such auxiliary functions are designed to be lightweight compared to classifiers so that they can significantly reduce the time cost. For example, Mao et al. [31] introduces a two-layer network that estimates the contrastive features, which can purify the adversarial images. Hwang et al. [17] propose AID-Purifier, an auxiliary discriminator based on information maximization principles that can transform adversarial images into natural images.
3 ZeroPur
3.1 Adversarial Purification in the Natural Image Manifold
Following the natural image manifold hypothesis [29, 44], natural images are assumed to reside on a specific manifold known as the natural image manifold. Given the classification loss function $\ell$, the learning process of DNNs finding optimal parameters $\theta$ can be formulated by:
$\min_{\theta}\ell(f\circ g(\bm{x}+\bm{\delta}^{*}),y;\theta),$ (1)
[…]
Adversarial attack crafts adversarial examples by optimizing the following objective:
$\max_{\|\bm{\delta}^{*}\|\leq\epsilon}\ell(f\circ g(\bm{x}+\bm{\delta}^{*}),y),$ (2)
where $\epsilon$ is the maximal norm which defines the set of allowed perturbations for a given example $\bm{x}$, in most contexts, such as the Projected Gradient Descent (PGD) [30], $\bm{\delta}^{*}$ is approximated by the local worst-cast $\bm{\delta}$.
We are starting with formulating the forward process of $f$. Suppose $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ is twice differentiable, then the forward of $f$ can be defined as:
$f_{1...l}(\bm{x})=f_{1}\circ f_{2}\circ...\circ f_{l}(\bm{x})=f_{l}(f_{l-1}(f_{2}...f_{1}(\bm{x}))),$ (3)
where $f_{i}$ typically refers to the output of the $\textit{i}^{th}$ layer in the function $f$. Based on the Taylor series, we have the following theorem:
Theorem 1.
Suppose $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ is twice differentiable at point $\bm{x}$. Let $e_{1}=\nabla f_{1}(\bm{x})^{T}\bm{\delta}+\frac{1}{2}\bm{\delta}^{T}\nabla^{2}f_{1}(\bar{\bm{x}})\bm{\delta}$, and there exists $w_{1}\in[0,1]$ such that $\bar{\bm{x}}=w_{1}\bm{x}+(1-w_{1})(\bm{x}+\bm{\delta})$, then we have the forward of $f$:
$f_{1...l}(\bm{x}+\bm{\delta})$ $=f_{1...l}(\bm{x})+e_{l}(\bm{\delta}),$ (4)
$where\ \ \ e_{l}(\bm{\delta})=\nabla f_{l}(f_{1...l-1}(\bm{x}))^{T}$ $e_{l-1}(\bm{\delta})+\frac{1}{2}e_{l-1}(\bm{\delta})^{T}\nabla^{2}\bar{f}_{1...l-1}(\bm{x})e_{l-1}(\bm{\delta}).$ (5)
where there exists $w_{l}\in[0,1]$ such that $\bar{f}_{1...l-1}(\bm{x})=w_{l}f_{1...l-1}(\bm{x})+(1-w_{l})(f_{1...l-1}(\bm{x})+e_{l-1}(\bm{\delta}))$.
The proof of Theorem 1 can be found in Appendix A.1. For notational simplicity, we denote $f_{1...l}(\bm{x})=f(\bm{x})$ and $e_{l}(\bm{\delta})=e(\bm{\delta})$ when $l$ is the last layer of $f$. Theorem 1 describes how the perturbation $\bm{\delta}$ is amplified layer by layer during the forward of $f$ so that the embedding of natural image $\bm{x}$ deviates manifold $\mathcal{M}$ and become the outlier, thereby demonstrates the reason for the incorrect decision made by $g$:
$g(f(\bm{x}+\bm{\delta}))=g(f(\bm{x})+e(\bm{\delta}))\neq g(f(\bm{x})),$ (6)
where $e(\bm{\delta})$ can be viewed as the deviation distance from the natural image manifold. As shown in Figure 2(a), the adversarial images induced by malicious attacks on natural samples seriously deviate from the natural image manifold.
By this formulation, adversarial purification can be naturally seen as the process of estimating $e({\bm{\delta}})$ in the embedding space, the approximation of the natural image manifold. This means the adversarial embedding $f(\bm{x}_{\mathrm{adv}})=f(\bm{x}+\bm{\delta})$ can return the manifold along the same path $-e(\bm{\delta})$. While precise estimation of $e(\bm{\delta})$ is challenging, it is fortunate that only an approximation, denoted as $\tilde{e}(\bm{\delta})$, is sufficient to move the adversarial embedding back within the manifold. As can be observed in Figure 2(e), although the adversarial image after the blur operation does not fall exactly onto the original image, this favorable direction is sufficient to serve as a proxy target to guide the adversarial image correctly onto the natural manifold, as shown in Figure 2(c)(d). We will elaborate on this novel purification pipeline ZeroPur in the following two Sections.
Figure 2: The impact of adversarial samples and their improvement through different operations, visualized using t-SNE. (a) AutoAttack [4] causes adversarial samples to deviate from the manifold; (b) Gaussian Blur, (c) Guided Shift and (d) ZeroPur (Ours) reverse this deviation to varying degrees. (e) Feature trajectories under adversarial attack and the blur operation, indicating that simple blurring operator can provide a promising direction toward the natural image manifold.
3.2 Guided Shift
Considering adversarial purification as an inverse process to adversarial attacks, we can emulate Eq.(2) to formulate the optimization objective used to estimate $\tilde{e}(\bm{\delta})$:
$\min_{\bm{\delta}_{\mathrm{pfy}}}\ell(f\ \circ$ $g(\bm{x}_{\mathrm{adv}}+\bm{\delta}_{\mathrm{pfy}}),y)$ (7)
$\mathrm{s.t.}\ \ \ \|\bm{\delta}_{\mathrm{pfy}}\|\leq\epsilon_{\mathrm{pfy}}$ $\mathrm{and}\ e(\bm{\delta}_{\mathrm{pfy}})=-\tilde{e}(\bm{\delta})\approx-e(\bm{\delta}),$ (8)
where $\bm{\delta}_{\mathrm{pfy}}$ and $\epsilon_{\mathrm{pfy}}$ are defined to correspond to $\bm{\delta}$ and $\epsilon$ in Eq.(2) to offset the perturbation. However, even if $\epsilon_{\mathrm{adv}}$ can be considered as a hyperparameter, the ground-truth label $y$ is not accessible. External model-based adversarial purification [56, 33, 25, 43] typically train a purification model to minimize the global $\ell_{\mathrm{pfy}}$. Auxiliary function-based adversarial method [31, 42, 17] tends to design a suitable $\ell_{\mathrm{pfy}}$ to complete purification without relying on external generative models. They need to retrain the classifier or parameterized auxiliary function to effectively cooperate in removing adversarial perturbations, which means these methods all introduce parameters $\Theta$ in $\ell_{\mathrm{pfy}}$:
$\min_{\bm{\delta}_{\mathrm{pfy}}}\ell_{\mathrm{pfy}}(f(\bm{x}_{\mathrm{adv}}+\bm{\delta}_{\mathrm{pfy}});\Theta).\$ (9)
Based on our assumption in Section 3.1 that adversarial images are outliers of the natural image manifold and the purification process can be considered as returning them to this manifold, we aim to design $\ell_{\mathrm{pfy}}$ without $\Theta$ by allowing them to shift towards the manifold adaptively:
$\min_{\bm{\delta}_{\mathrm{pfy}}}\ell^{\star}_{\mathrm{pfy}}(f(\bm{x}_{\mathrm{adv}}+\bm{\delta}_{\mathrm{pfy}}).$ (10)
We are starting to investigate whether a simple image transformation, such as color jitter, grayscale, Gaussian blur, solarization, and equalization, can shift adversarial images towards the natural image manifold. Fortunately, as shown in Figure 2(e), the blurring operation nearly provides a genuine direction toward the natural samples for the adversarial examples. This phenomenon suggests that a simple blurring operator can bring adversarial examples closer to the manifold.
Algorithm 1 Guided Shift
1:Adversarial example $\bm{x}_{\mathrm{adv}}$, iterations $T_{\mathrm{g}}$, step size $\eta_{1}$, a classifier $f$, purification bound $\epsilon_{\mathrm{pfy}}$, and the blurring operator $\mathrm{blur}(\cdot)$.
2:Guided result $\bm{x}_{\mathrm{g}}$
3:Random start $\bm{x}_{\mathrm{g}}^{0}\leftarrow\bm{x}_{\mathrm{adv}}+\varepsilon$
4:for $t=0,1,2,...,T_{\mathrm{g}}-1$ do
5:    $\bm{\tilde{x}}_{\mathrm{g}}^{t}\leftarrow\mathrm{blur}(\bm{x}_{\mathrm{g}}^{t})$
6:    $\bm{z}\leftarrow f(\bm{x}_{\mathrm{g}}^{t}),\bm{\tilde{z}}\leftarrow f(\bm{\tilde{x}}_{\mathrm{g}}^{t})$
7:    $\bm{x}_{\mathrm{g}}^{t+1}\leftarrow\bm{x}_{\mathrm{g}}^{t}+\eta_{1}\cdot\mathrm{sgn}\bigl{(}\nabla_{\bm{x}_{\mathrm{g}}^{t}}d(\bm{z},\bm{\tilde{z}})\bigr{)}$
8:    $\bm{x}_{\mathrm{g}}^{t+1}\leftarrow\mathrm{Clip}(\bm{x}^{t+1}_{\mathrm{g}},-\epsilon_{\mathrm{pfy}},\epsilon_{\mathrm{pfy}})$
9:    $\bm{x}_{\mathrm{g}}^{t+1}\leftarrow\mathrm{Clip}(\bm{x}^{t+1}_{\mathrm{g}},0,1)$
10:end for
11:$\bm{x}_{\mathrm{g}}\leftarrow\bm{x}_{\mathrm{g}}^{T_{\mathrm{g}}}$
Figure 3: Details of GS on CIFAR-10.
Therefore, we can accumulate this reversion to guide the adversarial examples back towards the manifold. For instance, we can pull the distance between adversarial examples and their blurred counterparts in the embedding space, which allows adversarial examples to converge closer to the natural image manifold. The distance of feature embeddings is defined by the Cosine Similarity:
$d(\bm{z}_{\mathrm{adv}},\bm{z}_{\mathrm{adv}}^{\prime})=\frac{\bm{z}_{\mathrm{adv}}\cdot\bm{z}_{\mathrm{adv}}^{\prime}}{\|\bm{z}_{\mathrm{adv}}\|\|\bm{z}_{\mathrm{adv}}^{\prime}\|},$ (11)
where $\bm{z}_{\mathrm{adv}}$ and $\bm{z}_{\mathrm{adv}}^{\prime}$ are the embeddings of adversarial examples and their blurred counterparts, respectively. we can shift the adversarial example $\bm{x}_{\mathrm{adv}}$ by the gradient of Eq.(11):
$\bm{x}_{\mathrm{adv}}^{+}=\bm{x}_{\mathrm{adv}}+\eta_{1}\cdot\nabla_{\bm{x}_{\mathrm{adv}}}d(f(\bm{x}_{\mathrm{adv}}),f(\bm{x}_{\mathrm{adv}}^{\prime})).$ (12)
Controlling the magnitude of the blurring applied carefully is essential to prevent excessive blurring that could render adversarial examples unrecognizable by the classifier. However, a small magnitude is not sufficient to shift images to reverse the distortion caused by adversarial perturbations. We therefore consider to move $\bm{x}_{\mathrm{adv}}$ by Eq.(12) with a small step size $\eta_{1}$ iteratively. This process is referred to as Guided Shift (GS). In each step of GS, we apply the same update rule:
$\bm{x}_{\mathrm{g}}^{t+1}=\bm{x}_{\mathrm{g}}^{t}+\eta_{1}$ $\cdot\nabla_{\bm{x}_{\mathrm{g}}^{t}}d(f(\bm{x}_{\mathrm{g}}^{t}),f(\bm{\tilde{x}}_{\mathrm{g}}^{t})),$ (13)
$\bm{\tilde{x}}_{\mathrm{g}}^{t+1}$ $=\mathrm{blur}(\bm{x}_{\mathrm{g}}^{t+1})$ (14)
where $\mathrm{blur}(\cdot)$ is the blurring operator and $\bm{x}_{\mathrm{g}}^{0}:=\bm{x}_{\mathrm{adv}}$. The workflow of GS is shown in Algorithm 1, where we use $\mathrm{sgn}(\cdot)$ to regulate the step size.
To demonstrate the process of GS, we employ two blurring operators: the Median filter and Gaussian blur to compute the cosine similarity between single-step result $\bm{x}_{\mathrm{g}}^{t}$ of GS and the natural image $\bm{x}$ ($\star$-$GS$-$N$), as well as between the blurred counterpart $\bm{\tilde{x}}_{\mathrm{g}}^{t}$ of $\bm{x}_{\mathrm{g}}^{t}$ and $\bm{x}$ ($\star$-$GS^{\prime}$-$N$), as shown in Figure 3. Interestingly, as GS progresses, both cosine similarities consistently increase, overcoming the inherent limitations of a single blur. Its effectiveness can be qualitatively described by Figure 2(c), where adversarial examples, guided by their blurred counterparts, gradually approach the natural manifold.
3.3 Adaptive Projection
We can also observe an interesting phenomenon in Figure 3. Regardless of the type of blurring operators, dash lines ($\star$-$GS^{\prime}$-$N$) and solid lines ($\star$-$GS$-$N$) tend to converge. Therefore, we can have the following assumption:
Assumption 1.
There is a limitation for Guided Shift in that the adversarial images can be guided by their blurred counterparts only toward the manifold rather than fully back to the manifold. In other words, this indicates that Guided Shift approximates a convex function.
Figure 4: Intuitive explanation of AP in the context of adversarial or natural $\bm{x}_{\mathrm{init}}$. PR: the perceptual regularization (Eq.(20)).
Algorithm 2 Adaptive Projection
1:Input image $\bm{x}_{\mathrm{init}}$, guided result $\bm{x}_{\mathrm{g}}$, iterations $T_{\mathrm{p}}$, a classifier $f$, purification bound $\epsilon_{\mathrm{pfy}}$ , and candidate layers $S$.
2:Adaptive result $\bm{x}_{\mathrm{p}}$
3:Initialize $\bm{x}_{\mathrm{p}}^{0}\leftarrow\bm{x}_{\mathrm{init}}$
4:Set step size $\eta_{2}\leftarrow\epsilon_{\mathrm{pfy}}/T_{\mathrm{p}}$
5:for $t=0,1,2,...,T_{\mathrm{p}}-1$ do
6:    for $l\in S$ do
7:        $\Delta\bm{u}_{\mathrm{g}}^{l}\leftarrow f_{l}(\bm{x}_{\mathrm{g}})-f_{l}(\bm{x}_{\mathrm{init}})$
8:        $\Delta\bm{u}_{\mathrm{p}}^{l}\leftarrow f_{l}(\bm{x}_{\mathrm{p}}^{t})-f_{l}(\bm{x}_{\mathrm{init}})$
9:        $\mathcal{L}\leftarrow\mathcal{L}+(-\Delta\bm{u}_{\mathrm{g}}^{l}\cdot\Delta\bm{u}_{\mathrm{p}}^{l})$
10:    end for
11:    $\mathcal{L}\leftarrow\frac{\lambda_{1}}{\|S\|}\mathcal{L}+\lambda_{2}\|\phi(\bm{x}_{\mathrm{p}}^{t}-\phi(\bm{x}_{\mathrm{init}}))\|_{2}$
12:    $\bm{x}_{\mathrm{p}}^{t+1}\leftarrow\bm{x}_{\mathrm{p}}^{t}-\eta_{2}\cdot\mathrm{sgn}(\nabla_{\bm{x}_{\mathrm{p}}^{t}}\mathcal{L})$
13:    $\bm{x}_{\mathrm{p}}^{t+1}\leftarrow\mathrm{Clip}(\bm{x}_{\mathrm{p}}^{t+1},-\epsilon_{\mathrm{pfy}},\epsilon_{\mathrm{pfy}})$
14:    $\bm{x}_{\mathrm{p}}^{t+1}\leftarrow\mathrm{Clip}(\bm{x}_{\mathrm{p}}^{t+1},0,1)$
15:end for
16:$\bm{x}_{\mathrm{p}}\leftarrow\bm{x}_{\mathrm{p}}^{T_{\mathrm{p}}}$
Certainly, this limitation arises from the exceedingly blurred images whose embeddings cannot be recognized correctly by the classifier. To fully return adversarial images to the manifold, we must overcome this limitation. A natural intuition is to allow current images to move adaptively instead of being guided by blurred counterparts, which requires us to target a basic direction for current images. Fortunately, the result of Guided Shift already provides this basic direction. Therefore, we propose Adaptive Projection (AP) that can be defined as:
$\max_{\bm{x}_{\mathrm{p}}}\lambda_{1}\frac{1}{\|S\|}\sum_{l\in S}\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})-\lambda_{2}\|\phi(\bm{x}_{\mathrm{p}})-\phi(\bm{x}_{\mathrm{init}})\|_{2}\ \ \ \ \mathrm{s.t.}\ \ \ \|\bm{x}_{\mathrm{p}}-\bm{x}_{\mathrm{init}}\|\leq\bm{\delta}_{\mathrm{pfy}},$ (15)
where $S\subseteq L$ as a candidate in the set $L=\{l_{1},l_{2},...,l_{m}\}$ of a $m$ layers model $f$, $\|S\|$ denotes the number of elements of the set $S$. $\bm{x}_{\mathrm{init}}$ represents the input image (i.e., $\bm{x}_{\mathrm{init}}=\bm{x}_{\mathrm{adv}}$ when the input image is adversarial). For the $l^{th}$ layer of classifiers, the first term of Eq. (15) can be written as:
$\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})=\Delta\bm{u}^{l}_{\mathrm{p}}\cdot\Delta\bm{u}^{l}_{\mathrm{g}},$ (16)
where $\bm{x}_{\mathrm{p}}$ is the final result, $\Delta\bm{u}^{l}_{\mathrm{p}}$ and $\Delta\bm{u}^{l}_{\mathrm{g}}$ are two vectors of flattened feature maps defined as follow:
$\Delta\bm{u}^{l}_{\mathrm{p}}=f_{l}(\bm{x}_{\mathrm{p}})-f_{l}(\bm{x}_{\mathrm{init}}),\ \ \ \Delta\bm{u}^{l}_{\mathrm{g}}=f_{l}(\bm{x}_{\mathrm{g}})-f_{l}(\bm{x}_{\mathrm{init}}),$ (17)
where $f_{l}$ denotes feature maps at layer $l$ of the classifier, $\bm{x}_{\mathrm{p}}$ is initialized by $\bm{x}_{\mathrm{init}}$. Maximizing Eq.(16) is equivalent to maximizing the projection of $\bm{u}^{l}_{\mathrm{p}}$ onto $\bm{u}^{l}_{\mathrm{g}}$ since $\|\bm{u}^{l}_{\mathrm{g}}\|$ is a constant. The increase in projection allows $\bm{x}_{\mathrm{p}}$ to move freely along the direction of GS, enabling it to independently approach the natural image manifold.
However, We typically do not know whether an image is adversarial in real-world applications (i.e., $\bm{x}_{\mathrm{init}}=\bm{x}$ when $\bm{x}$ is natural), repeatedly guiding images by blurred counterparts may cause natural images to deviate from the manifold. As shown in Figure 4, this deviation will be enhanced by AP. Therefore, we introduce a perceptual regularization term in Eq.(15). Let $F_{1}$ is the dynamic of Eq.(16):
$F_{1}=\lambda_{1}\cdot\nabla_{\bm{x}_{\mathrm{p}}}\frac{1}{\|S\|}\sum_{l\in S}\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})\ ,$ (18)
If there exists a momentum $F_{2}$, such that:
$\left\{\begin{array}[]{lll}F_{1}>F_{2},&\bm{x}_{\mathrm{init}}\ is\ adversarial,\\ F_{1}<F_{2}&\bm{x}_{\mathrm{init}}\ is\ natural,\\ \end{array}\right.$ (19)
then maximizing Eq. (16) allows the result $\bm{x}_{p}$ to locate in the natural image manifold. In this work, the dynamic $F_{2}$ can be computed by LPIPS distance [60]. The LPIPS distance $d(\bm{x}_{1},\bm{x}_{2})$ between images $\bm{x}_{1}$ and $\bm{x}_{2}$ is then defined as $d(\bm{x}_{1},\bm{x}_{2})\triangleq\|\phi(\bm{x}_{1})-\phi(\bm{x}_{2})\|_{2}$. Let $\hat{f}(\bm{x})$ denote channel-normalized activations at the $l$-th layer of the classifier. Then, $\phi(\bm{x})\triangleq(\frac{\hat{f}(\bm{x})}{\sqrt{w_{1}h_{1}}},...,\frac{\hat{f}(\bm{x})}{\sqrt{w_{m}h_{m}}})$, where $w_{l}$ and $h_{l}$ are the width and height of activations of layer $l$, respectively.
Finally, the momentum $F_{2}$ can be written as:
$\centering F_{2}=-\lambda_{2}\cdot\nabla_{\bm{x}_{\mathrm{p}}}\|\phi(\bm{x}_{\mathrm{p}})-\phi(\bm{x}_{\mathrm{init}})\|_{2}.\@add@centering$ (20)
In contrast to the original LPIPS implementation, we do not require this distance to approximate human perceptual judgments. Instead, we aim for the perceptual distance between $\bm{x}_{\mathrm{g}}$ and $\bm{x}_{\mathrm{init}}$ based on the classifier itself to be small. Hence, we can directly use $d(\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{init}})$ to evaluate the distance from the perspective of the classifier. The workflow of AP is shown in Algorithm 2.
4 Experiments
Datasets and Metrics. Three benchmarks CIFAR-10 [21], CIFAR-100 [21], and ImageNet-1K [8] are considered to evaluate our method. We compare our method with the state-of-the-art adversarial training methods reported in standard benchmark RobustBench [5], as well as with other representative adversarial training and purification methods [42, 31, 26, 6, 22, 46, 15, 56, 48, 33, 25, 27, 43]. In all experiments, we consider two metrics to evaluate the performance of all methods: standard accuracy (SA) and robust accuracy (RA). The standard accuracy measures the performance of the defense method on natural images, while the robust accuracy measures the performance on adversarial images.
Adversarial attacks and victim classifiers. We evaluate our methods with three attacks: AutoAttack [4], DI2-FGSM [54], and BPDA [2]. AutoAttack is a powerful adaptive attack commonly used in most defense studies. We also use DI2-FGSM to seek whether an attack robust to the blurring operator can affect our method. Additionally, we consider BPDA to attack our purification module in the worst-case scenario. These attacks will attack three victim classifiers, including ResNet-18 [13], ResNet-50 [13], and WideResNet-28-10 [58]. For more information about these attacks, refer to Appendix A.3.
Table 1: Comparison with AT &\&& ABP methods under AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$). $\dagger$ indicates this method uses WideResNet-34-10 as a classifier.
Type Require Training ResNet-18 WideResNet-28-10
Cls Aux Method SA(%) RA(%) Method SA(%) RA(%)
CIFAR-10
AT (Gowal et al., 2021) 87.35 59.12 (Gowal et al., 2021) 87.50 63.99
(Sehwag et al., 2021) 84.59 56.19 (Pang et al., 2022) 88.61 61.40
(Rade et al., 2021) 89.02 58.17 (Xu et al., 2023) 93.69 65.62
(Addepalli et al., 2022) 85.71 52.90 (Wang et al., 2023b) 92.44 67.31
(Li and Spratling, 2023) 83.45 52.52 (Cui et al., 2024) 92.16 67.73
[…]
(Mao et al., 2021) - 58.20 (Mao et al., 2021) - 67.15
(Kulkarni and Weng, 2024) 85.18 54.53 (Hwang et al., 2023$\dagger$ 82.22 56.63
GS 49.56 58.82 GS 50.24 59.68
GS+AP (ZeroPur) 92.56 69.62 GS+AP (ZeroPur) 91.81 68.60
CIFAR-100
AT (Rade et al., 2021) 61.50 29.50 (Pang et al., 2022) 63.66 31.67
(Addepalli et al., 2022) 65.45 28.58 (Rebuffi et al., 2021) 62.41 33.03
[…]
(Mao et al., 2021) - 25.45 (Mao et al., 2021) - 33.16
(Kulkarni and Weng, 2024) 65.37 32.55 (Hwang et al., 2023$\dagger$ 64.73 32.86
GS 30.42 34.35 GS 26.56 32.06
GS+AP (ZeroPur) 52.45 41.42 GS+AP (ZeroPur) 64.88 34.58
4.1 Quantitative Evaluation and Comparison
For adversarial defense, adversarial training (AT) involves adversarial examples in the classifier training, and auxiliary-based purification (ABP) introduces an auxiliary function to cooperate with the victim classifier. These methods do not rely on external generative models but still necessitate training auxiliary functions or retraining classifiers. External model-based purification (EBP) fine-tunes or retrains an external generative model to remove adversarial perturbations in adversarial images. Since the performance of EBP is ensured by the modeling capability of generative models, it typically outperforms ABP. In this section, for a fair comparison, we will first compare our method with both AT and ABP and then compare it with EBP. In Table 1,  3, and  3, the terms ’Ext’, ’Aux’, and ’Cls’ under ’Require Training’ indicate whether the methods require retraining external generative models, auxiliary functions, or victim classifiers.
Comparison with AT &\&& ABP. Table 1 reports the standard and robust accuracy against AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$). ZeroPur achieves consistent and significant improvements in robust accuracy, particularly on ResNet-18, with robust accuracies of $69.62\%$ and $41.42\%$ on CIFAR-10 and CIFAR-100, respectively. Additionally, we report the comparison on ImageNet-1K in Table 3, where ZeroPur improves robust accuracy by $30.06\%$ compared to AT and ABP methods, representing a substantial gain. These results demonstrate that ZeroPur effectively guides adversarial images toward the natural image manifold. Notably, our method performs better on smaller architectures like ResNet-18, likely because larger models have more complex feature spaces, making them more sensitive to attacks, which increases the difficulty of mapping adversarial images to natural ones using our method.
Comparison with EBP. Although generative model-based purification methods achieve outstanding performance due to their large scale and high computation, our method still achieves state-of-the-art or comparable performance. Table 3 and Table 3 report the comparison result between ZeroPur and EBP. For CIFAR-10, ZeroPur achieves robust accuracy comparable to existing EBP defense methods, and a slightly tweaked ZeroPur (see the first part of Section 4.2) can achieve state-of-the-art performance over EBP methods. On ImageNet, our method achieves SOTA robustness, but with a slight decrease in natural accuracy. For example, even compared to the latest Mimicdiffusion [43], our method demonstrates a $5.26\%$ improvement in robust accuracy with only a $3.15\%$ loss in clean accuracy. Figure 6 shows the visualizations and differences resulting from applying the two-step process of ZeroPur to adversarial examples. The reduced difference between the purified and natural images, along with the increased difference from adversarial images, demonstrates the effectiveness of our method. More experimental results and visualizations (including failure cases) are provided in Appendix A.4.
Table 2: Comparison with EBP methods on CIFAR-10 under AutoAttack $\ell_{\infty}(\epsilon=8/255)$.
Require Training WideResNet-28-10
Cls Ext Method SA(%) RA(%)
(Sun et al., 2019) 82.22 67.92
(Hill et al., 2020) 84.12 78.91
(Yoon et al., 2021) 86.14 80.24
(Ughini et al., 2022) - 59.57
(Nie et al., 2022) 89.02 70.64
(Lee and Kim, 2023) 90.16 70.47
(Lin et al., 2024) 90.62 72.85
[…]
ZeroPur 85.02 82.76
Table 3: Comparison on ImageNet-1K under AutoAttack $\ell_{\infty}(\epsilon=4/255)$.
Training ResNet-50
Cls Aux Ext Method SA(%) RA(%)
Comparison with AT &\&& ABP
(Salman et al., 2020) 64.02 34.96
(Wong et al., 2020) 55.62 26.24
(Bai et al., 2021) 67.38 35.51
[…]
(Kulkarni and Weng, 2024) 64.10 37.36
ZeroPur 63.77 67.42
Comparison with EBP
(Nie et al., 2022) 67.79 40.93
(Lee and Kim, 2023) 70.74 46.31
(Song et al., 2024) 66.92 62.16
[…]
4.2 Discussion on ZeroPur
Evaluating ZeroPur on protecting the victim classifier trained with different data augmentation. To further discuss the performance of ZeroPur, we consider three data augmentation strategies (’Vanilla’, ‘Base’, ‘Strong’) to train victim classifiers (See Appendix A.2 for details). ‘Base’ is the most common data augmentation in training classifiers. In the experiment, ZeroPur using a Median filter ($3\times 3$ window size) is applied in ’Vanilla’ classifiers and Gaussian blur ($\sigma=1.2$) in ‘Base’ and ’Strong’ classifiers, please refer to Table 6 in Appendix A.4. Figure 5(a) and Figure 5(b) report the standard and robust accuracy of ZeroPur under three different training recipes for ResNet-18 and WideResNet-28-10, respectively. It can be observed that the ’Strong’ data augmentation benefits our method. The reason for this is that models trained with such augmentation exhibit better perceptual capability for blurred images, enabling them to more effectively guide adversarial images towards the natural image manifold and facilitate precise adaptive projection. The performance of ZeroPur in Table 3 also demonstrates this phenomenon.
(a) ResNet-18
(b) WideResNet-28-10
(c) Ablation of PR
(d) Blur vs. TVM
Figure 5: SA and RA (%) of classifiers trained with three-level data augmentation on (a) ResNet-18 and (b) WideResNet-28-10. (c) Ablation Study of perceptual regularization on CIFAR-10. (d) Robust accuracy against DI2-FGSM and AutoAttack on CIFAR-10 by Blur and TVM.
Impact of perceptual regularization. Figure 5(c) illustrates the impact of perceptual regularization in Eq.(15) on standard accuracy. The comparison between models with and without PR (red vs. green) reveals that incorporating PR leads to a notable improvement in standard accuracy with almost no loss in robust accuracy. This trend is consistent across different architectures, ResNet-18 (R18) and WideResNet-28-10 (WRN), confirming the effectiveness of PR in the adaptive projection stage.
Figure 6: Visual examples of ZeroPur against $\ell_{\infty}$ threat model ($\epsilon=4/255$) on ImageNet-1K. The red denotes error prediction and the green denotes correct prediction. $\Delta_{\mathrm{A}\text{-}\mathrm{B}}$ represents the normalized absolute difference between A and B.
ZeroPur defense against adversarial attacks robust to blurring. ZeroPur employs blurring to project adversarial images back onto the natural image manifold. Although certain attacks, such as DI2-FGSM[54], exhibit inherent resistance to blurring, they are still not entirely immune to ZeroPur. As shown in Figure 5(d), replacing the blurring operator with TVM [12] leads to improved purification. The robust accuracy against AutoAttack increases to $75.96\%$ without training, which outperforms the optimal performance of GS(Blur)+AP by $6.34\%$. This result also demonstrates that ZeroPur is not limited to blurring and can generalize to other operators.
[…]
Model PGD-10 PGD-20 PGD-40
Vanilla WRN 48.63 47.00 45.75
R18 51.30 51.30 51.10
Base WRN 34.16 33.38 34.42
R18 39.88 39.90 39.48
Strong WRN 68.84 69.45 68.50
R18 70.65 70.45 70.43
ZeroPur defense against attacks that bypass the purification module. The efficiency of purification methods can be challenged by the BPDA attack that approximates gradients, bypassing the purification module. Table 4.2 reports the performance of ZeroPur defense against BPDA attack. We can observe that the robust accuracy of ZeroPur decreases, but remains stable, indicating that ZeroPur provides acceptable defense. As mentioned earlier, the larger model WRN consistently exhibits slightly worse robustness against BPDA compared to the smaller model R18, while achieving better natural accuracy than R18 (See Table 5 in Appendix A.2).
5 Conclusion
We propose a succinct training-free method for adversarial purification, named ZeroPur. Our method significantly outperforms previous state-of-the-art adversarial training and auxiliary-based purification methods, while demonstrating comparability with external model-based purification methods. Despite the improvements, ZeroPur has a major limitation: the absence of external models restricts our ability to enhance defense against strong adaptive attacks. To overcome this challenge, it is necessary to design a module that emulates recent works on purification by diffusion models capable of effectively defending BPDA.
Appendix A Appendix
A.1 Proofs
Theorem 1.
Suppose $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ is twice differentiable at point $\bm{x}$. Let $e_{1}=\nabla f_{1}(\bm{x})^{T}\bm{\delta}+\frac{1}{2}\bm{\delta}^{T}\nabla^{2}f_{1}(\bar{\bm{x}})\bm{\delta}$, and there exists $w_{1}\in[0,1]$ such that $\bar{\bm{x}}=w_{1}\bm{x}+(1-w_{1})(\bm{x}+\bm{\delta})$, then we have the forward of $f$:
$f_{1...l}(\bm{x}+\bm{\delta})$ $=f_{1...l}(\bm{x})+e_{l}(\bm{\delta}),$ (1)
$where\ \ \ e_{l}(\bm{\delta})=\nabla f_{l}(f_{1...l-1}(\bm{x}))^{T}$ $e_{l-1}(\bm{\delta})+\frac{1}{2}e_{l-1}(\bm{\delta})^{T}\nabla^{2}\bar{f}_{1...l-1}(\bm{x})e_{l-1}(\bm{\delta}).$ (2)
where there exists $w_{l}\in[0,1]$ such that $\bar{f}_{1...l-1}(\bm{x})=w_{l}f_{1...l-1}(\bm{x})+(1-w_{l})(f_{1...l-1}(\bm{x})+e_{l-1}(\bm{\delta}))$.
Proof.
By a first-order Taylor series, we have:
$f(\bm{x})=f(\bm{x}_{0})+\nabla f(\bm{x}_{0})^{T}(\bm{x}-\bm{x}_{0})+\frac{1}{2}(\bm{x}-\bm{x}_{0})^{T}\nabla^{2}f(\bar{\bm{x}})(\bm{x}-\bm{x}_{0}).$ (3)
There exists $w\in[0,1]$ such that $\bar{\bm{x}}=w\bm{x}+(1-w)\bm{x}_{0}$. for the forward of an adversarial example $\bm{x}+\bm{\delta}$ in the first layer $f_{1}$, we have:
$f_{1}(\bm{x}+\bm{\delta})=f_{1}(\bm{x})+\underbrace{{\nabla f_{1}(\bm{x})^{T}\bm{\delta}+\frac{1}{2}\delta^{T}\nabla^{2}f_{1}(\bar{\bm{x}})\bm{\delta}}}_{e_{1}(\bm{\delta})}.$ (4)
Plugging $f_{2}$ into Eq.(4), we also have:
$f_{1,2}(\bm{x}+\bm{\delta})$ $=f_{2}(f_{1}(\bm{x}+\bm{\delta}))=f_{2}(f_{1}(\bm{x})+e_{1}(\bm{\delta}))$ (5)
$=\underbrace{f_{2}(f_{1}(\bm{x}))}_{f_{1,2}(\bm{x})}+\underbrace{\nabla f_{2}(f_{1}(\bm{x}))^{T}e_{1}(\bm{\delta})+\frac{1}{2}e_{1}(\bm{\delta})^{T}\nabla^{2}\bar{f}_{1}(\bm{x})e_{1}(\delta)}_{e_{2}(\bm{\delta})}.$ (6)
when $\bm{x}+\bm{\delta}$ is forward through the $l$-th layer, we have Theorem 1. $\square$
A.2 Implementation Details of Classifiers
Our experiments on CIFAR-10 and CIFAR-100 included ResNet-18 and WideResNet-28-10 trained on three different data augmentation strategies: ‘Vanilla’, ‘Base’, and ‘Strong’. Table 5 shows the training settings and natural accuracy. ‘Base’ is the most common data augmentation used by adversarial training methods and adversarial purification methods for better natural accuracy.
Table 5: The adopted data augmentation and natural accuracy (%) of victim classifiers. The enumerated data augmentation are, in order, ResizeCrop, ColorJitter, Grayscale, Solarization, Equalization, and HorizontalFlip.
Victim Classifier Case Trained On Natural Accuracy (%)
ReCrop. ColorJ. GrayS. GauBlur. Solar. Equal. HorFlip. CIFAR-10 CIFAR-100
ResNet-18 Vanilla 83.80 57.01
Base 93.10 71.58
Strong 91.08 68.53
WRN-28-10 Vanilla 91.34 71.57
Base 93.83 74.59
Strong 91.10 67.65
All classifiers were trained with the SGD optimizer with a cosine decay learning rate schedule and a linear warm-up period of 10 epochs. The weight decay is $5.0\times 10^{-4}$ and the momentum is $0.9$. The initial learning rate is set to $0.1$. Classifiers were trained for 120 epochs on 4 Tesla V100 GPUs, where the batch size is 512 per GPU for ResNet-18 and 128 per GPU for WideResNet-28-10.
A.3 Implementation Details of Adversarial Attacks
AutoAttack [4].  We use AutoAttack to compare with the start-of-the-art methods. The robust classifier for adversarial training methods provided by RobustBench [5] benchmark is available at https://robustbench.github.io. The code for adversarial purification methods is provided by their respective papers.
There are two versions of AutoAttack: (i) the Standard including AGPD-CE, AGPD-T, FAB-T, and Square, and (ii) the Rand version including APGD-CE and APGD-DLR. Considering that most of the adversarial purifications choose the Rand version, all the performance in this work we report is also in the Rand version. Code is available at https://github.com/fra31/auto-attack.
DI2-FGSM [54]. DI2-FGSM crafts adversarial examples by applying various transformations, enhancing their robustness against blurring operations. We use this attack to evaluate ZeroPur in Section 4.2, implemented by torchattacks [20]. Code is available at https://github.com/Harry24k/adversarial-attacks-pytorch.
BPDA [2]. We use BPDA, approximating the gradient of the purifier module (ZeroPur) as 1 during the backward pass. The 10, 20, and 40 iterations are applied in our experiment. Other settings are the same as those used in PGD Attack.
Table 6: SA / RA (%) on CIFAR-10 with different augmentations under varying blurring operators. The bold indicates the best harmonic mean of SA and RA, measured by the F1 Score.
Operator Level ResNet-18 WRN-28-10
Vanilla Base Strong Vanilla Base Strong
Median $3\times 3$ 82.04/ 59.38 92.34/ 72.27 90.90/ 60.74 91.22/ 59.16 91.81/ 67.56 90.65/ 70.01
$5\times 5$ 83.73/ 34.17 91.87/ 64.81 90.92/ 44.63 91.06/ 39.81 91.34/ 58.41 88.97/ 50.49
$7\times 7$ 83.66/ 27.25 91.56/ 54.41 90.87/ 34.68 90.97/ 28.94 90.34/ 48.26 87.70/ 40.58
Gaussian $\sigma=0.6$ 83.57/ 38.82 91.64/ 65.43 90.72/ 64.92 91.09/ 53.21 90.02/ 58.31 90.68/ 73.79
$\sigma=1.2$ 83.63/ 26.20 92.56/ 69.62 90.05/ 83.84 91.14/ 27.97 91.81/ 68.80 85.02/ 82.76
$\sigma=1.8$ 83.69/ 20.45 92.36/ 60.67 90.79/ 67.89 51.26/ 26.82 35.95/ 59.91 80.27/ 78.82
A.4 More Experimental Results
In Table 7 and Table 8, we compare our method with AT, ABP, and EBP methods on CIFAR-10 against $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$ threat model, detailing training requirements for each method. We also report the performance of ZeroPur when victim classifiers use strong data augmentation (ZeroPur). Among all methods, ZeroPur generally achieves comparable standard and robust accuracy to any other method, while its enhanced version ZeroPur consistently attains the best robust accuracy.
To evaluate the effect of training strategies and blur types on ZeroPur, Table 6 presents its standard and robust accuracy under different augmentations (Vanilla, Base, Strong) using Median filters ($3\times 3$, $5\times 5$, $7\times 7$) and Gaussian blurs ($\sigma=0.6$, $1.2$, $1.8$). The results show that if the victim classifiers are not trained with strong data augmentation, the low-quality embeddings induced by stronger blur operators will hinder ZeroPur’s ability to guide adversarial images toward the natural manifold. In contrast, victim classifiers trained with strong data augmentation inherently possess the ability to recognize blurred images, enabling ZeroPur to benefit from high-intensity blurring. For commonly used base classifiers, they generally perform well under moderate blur intensity.
We provide additional visualization examples in Figure 7 to further validate the effectiveness and interpretability of ZeroPur. The progression from adversarial images to GS-purified images and then to AP-purified images shows decreasing differences from natural images (gradually transitioning from yellow to blue), indicating the effectiveness of our method in purification. We observe that in cases where GS successfully purifies adversarial samples, the difference between GS-purified and natural images (denoted as $\Delta_{\text{GS-Nat.}}$) displays scattered, non-semantic noise. In contrast, for cases where GS fails to purify, the difference $\Delta_{\text{GS-Nat.}}$ still shows semantic perturbations. Furthermore, the similarity between $\Delta_{\text{AP-Adv.}}$ and $\Delta_{\text{Adv.-Nat.}}$ suggests that our method (GS+AP, i.e., ZeroPur) focuses purification mainly on regions influenced by adversarial perturbations. However, there are still some failure cases. It is not difficult to observe that the inconsistency between the predictions of Adv. and GS already appears at the beginning. This highlights the importance of the initial purification direction, which may be the key to ZeroPur’s success.
Table 7: Comparison with three types methods on CIFAR-10 under AutoAttack $\ell_{\infty}(\epsilon=8/255)$.
Defense Type Method Training Architecture SA(%) RA(%)
AT (Gowal et al., 2021) Classifier ResNet-18 87.35 59.12
(Gowal et al., 2021) Classifier WRN-28-10 87.50 63.99
(Sehwag et al., 2021) Classifier ResNet-18 84.59 56.19
(Rade et al., 2021) Classifier ResNet-18 89.02 58.17
(Addepalli et al., 2022) Classifier ResNet-18 85.71 52.90
(Pang et al., 2022) Classifier WRN-28-10 88.61 61.40
(Xu et al., 2023) Classifier WRN-28-10 93.69 65.62
(Wang et al., 2023b) Classifier WRN-28-10 92.44 67.31
(Li and Spratling, 2023) Classifier ResNet-18 83.45 52.52
(Cui et al., 2024) Classifier WRN-28-10 92.16 67.73
[…]
(Hill et al., 2020) EBM+LD WRN-28-10 84.12 78.91
(Yoon et al., 2021) DSM+LD WRN-28-10 86.14 80.24
(Ughini et al., 2022) DeepFill [57] WRN-28-10 - 59.57
(Nie et al., 2022) DDPM [16] WRN-28-10 89.02 70.64
(Nie et al., 2022) DDPM [16] WRN-70-16 90.07 71.29
(Lee and Kim, 2023) DDPM [16] WRN-28-10 90.16 70.47
(Lee and Kim, 2023) DDPM [16] WRN-70-16 90.53 70.31
(Lin et al., 2024) MAE [14] WRN-28-10 90.62 72.85
(Lin et al., 2024) MAE [14] WRN-70-16 91.99 76.37
[…]
(Shi et al., 2021) Classifier WRN-28-10 91.89 68.56
(Mao et al., 2021) Auxiliary branch ResNet-18 - 58.20
(Mao et al., 2021) Auxiliary branch WRN-28-10 - 67.15
(Hwang et al., 2023) Auxiliary branch WRN-34-10 82.22 56.63
ZeroPur N/A ResNet-18 92.56 69.62
ZeroPur N/A ResNet-18 90.05 83.84
ZeroPur N/A WRN-28-10 91.81 68.60
ZeroPur N/A WRN-28-10 85.02 82.76
Table 8: Comparison with three types methods on CIFAR-10 under AutoAttack $\ell_{2}(\epsilon=0.5)$.
Defense Type Method Training Architecture SA(%) RA(%)
AT (Rebuffi et al., 2021) Classifier ResNet-18 90.33 75.86
(Sehwag et al., 2021) Classifier ResNet-18 89.76 74.41
(Wang et al., 2023b) Classifier WRN-28-10 95.16 83.68
(Rebuffi et al., 2021) Classifier WRN-28-10 91.79 78.80
(Sehwag et al., 2021) Classifier WRN-34-10 90.93 77.24
[…]
(Ughini et al., 2022) DeepFill [57] WRN-28-10 - 45.12
(Nie et al., 2022) DDPM [16] WRN-28-10 91.03 78.58
(Nie et al., 2022) DDPM [16] WRN-70-16 92.68 80.60
(Lin et al., 2024) MAE [14] WRN-28-10 90.62 80.47
(Lin et al., 2024) MAE [14] WRN-70-16 91.99 81.35
[…]
ZeroPur N/A ResNet-18 90.77 86.56
ZeroPur N/A WRN-28-10 72.38 72.12
ZeroPur N/A WRN-28-10 81.52 86.61
Figure 7: Visual examples of ZeroPur against $\ell_{\infty}$ threat model ($\epsilon=4/255$) on ImageNet-1K. The red denotes error prediction and the green denotes correct prediction. $\Delta_{\mathrm{A}\text{-}\mathrm{B}}$ represents the normalized absolute difference between A and B. (Zoom in for a clearer view.)